# A new approach to principal–agent problems with volatility control

### v1

@@ 5 Conclusion
In this paper, we develop a novel, simpler approach to solving principal–agent problems with volatility control. Specifically, while the existing approach, introduced by Cvitanić, Possamaï, and Touzi (2018) [9], relies on the theory of second-order BSDEs, the approach we propose allows us to solve this type of problem solely using BSDEs. More precisely, in [9], the main and challenging result is to prove the equivalence between the original principal’s problem, $V_{\rm P}$ in 1, and its reformulation as a more standard stochastic control problem, $\widetilde{V}_{\rm P}$ in 2. While it is clear from the restriction to contracts of the form (2.15) that $V_{\rm P}\geq\widetilde{V}_{\rm P}$, the converse inequality requires to prove that such restriction is without loss of generality, which necessitates the use of second-order BSDEs. Even though this approach indeed allows a thorough analysis of the principal–agent problem with volatility control defined by [9], and therefore of the applications arising from this framework, the mathematical complexity of this approach hampers the study of natural and relevant extensions, such as to multi-agent frameworks. Moreover, in our opinion, this mathematically challenging approach also prevents a wider dissemination of these models and their relevant applications in many fields, from operation research to economics.
@@
In contrast, we introduce in this paper an alternative ‘first-best’ formulation in 3, in which the principal directly controls the density of the output’s quadratic variation. This problem is purely theoretical: it can be seen as a dual to 1, and is not necessarily meaningful in terms of application. However, compared to 1, 3 can be solved in a more straightforward way, following the so-called Sannikov’s trick, as introduced by Sannikov (2008) [29]. More precisely, one can prove using BSDEs that the restriction to contracts of the form (3.12) is without loss of generality, which implies that the principal’s problem in this ‘first-best’ reformulation is actually equivalent to a more standard stochastic control problem, namely $\widetilde{V}_{\rm P}^{\circ}$ defined in 4. It then remains to show that the value $V_{\rm P}^{\circ}$ of the ‘first-best’ reformulated problem, which naturally satisfies $V_{\rm P}^{\circ}\geq V_{\rm P}$, can be achieved. For this, it suffices to use the contract form (2.11) considered in [9], and thus show that 2 and 4 are actually equivalent. In other words, instead of choosing the sensitivity parameter $\Gamma$ in the contract form (2.11), one can directly optimise the density process $\Sigma$ of the output’s quadratic variation. Conversely, given a density process $\Sigma$, one can recover from the one-to-one correspondence in Lemma 3.11 the appropriate pay-for-performance parameter $\Gamma$ in the contract form (2.11). We finally obtain that the problems introduced throughout this paper are all equivalent, which is graphically summarised by Figure 1.
1234Lemma 2.5, or [9]Lemma 3.4Theorem 3.9By 2BSDEs in [9]Theorem 3.10Lemma 3.11
@@
We nevertheless wish to mention that, at the end, and depending on the framework under consideration, 2 may still be easier to solve than 4. Indeed, as illustrated through Section 4, the agent’s problem 4 involves a constrained optimisation, making it potentially more complex to solve in practice. However, this does not undermine the effectiveness of our approach, since its main goal is to provide a theoretically simpler method to establish the optimality of the relevant contract forms. Ultimately, since all problems are equivalent, this approach offers the flexibility to choose the most practical problem to solve. We also wish to emphasise once more that the contract form (2.11) introduced in [9] is essential for proving the equivalence between the original problem and its ‘first-best’ reformulation. Nevertheless, the key message of our approach is that we do not need to rely on 2BSDEs to demonstrate the optimality of these contracts; their optimality is ensured by the fact that the ‘first-best’ value is achieved.
We believe that this alternative, simpler approach will provide a more straightforward pathway for extending principal–agent problems to more complex frameworks, involving for example multi-agent settings, mean-field interactions, or non-continuous output processes. For instance, solving principal–multi-agent problems with volatility control via the original approach in [9] requires the consideration of ‘multidimensional 2BSDEs’, as highlighted in [18], for which the theory is actually not yet developed. In contrast, following our proposed approach, one could rely on the more mature theory of multidimensional BSDEs to show the equivalence between the original problem and its ‘first-best’ reformulation, thereby ensuring optimality of appropriate revealing contracts without relying on multidimensional 2BSDEs. Furthermore, this simplified approach shall facilitate practical and more widespread implementation of principal–agent problems with volatility control in diverse fields, beyond mathematical finance. Indeed, such models are already used notably for energy-related applications, see [1, 11, 2], and may also prove useful in other domains.

### v2

@@ 5 Conclusion
Principal–agent problems with volatility control are notoriously challenging to study. Specifically, the only existing approach so far, introduced by Cvitanić, Possamaï, and Touzi (2018) [9], relies on the sophisticated theory of second-order BSDEs. More precisely, the authors establish the equality between the principal’s value $V_{\rm P}$ in the original problem (Problem 1) and the value $\widetilde{V}_{\rm P}$ of a more standard stochastic control problem (Problem 2). While it is clear from the restriction to contracts of the form (2.15) that $V_{\rm P}\geq\widetilde{V}_{\rm P}$, the converse inequality requires to prove that such restriction is without loss of generality, which necessitates the use of second-order BSDEs. Even though this approach indeed allows a thorough analysis of the principal–agent problem with volatility control, and therefore of the applications arising from this framework, the mathematical complexity of this approach hampers the study of natural and relevant extensions, especially to multi-agent frameworks. Moreover, this mathematically challenging approach also prevents a wider dissemination of these models and their relevant applications in many fields, from operation research to economics.
@@
In contrast, the novel approach we develop only relies on the theory of BSDEs, thanks to the introduction of a ‘first-best’ counterpart of the original problem, in which the principal directly controls the density of the output’s quadratic variation (Problem 3). This alternative problem is purely theoretical: it can be seen as a dual to Problem 1, and is not necessarily meaningful in terms of application. Nevertheless, compared to Problem 1, Problem 3 can be solved in a more straightforward way, following the so-called Sannikov’s trick. More precisely, one can prove using BSDEs that the restriction to contracts of the form (3.13) is without loss of generality, which implies that the principal’s problem in this ‘first-best’ reformulation is actually equivalent to another standard stochastic control problem, namely $\widetilde{V}_{\rm P}^{\circ}$ defined in Problem 4. It then remains to show that the value $V_{\rm P}^{\circ}$ of the ‘first-best’ reformulated problem, which naturally satisfies $V_{\rm P}^{\circ}\geq V_{\rm P}$, can be achieved. For this, it suffices to use the contract form (2.11) considered in [9], and thus show that Problem 2 and Problem 4 are actually equivalent. As a result, instead of choosing the sensitivity parameter $\Gamma$ in the contract form (2.11), one can directly optimise the density process $\Sigma$ of the output’s quadratic variation. Conversely, given a density process $\Sigma$, one can recover from the correspondence in Lemma 3.13 the appropriate pay-for-performance sensitivity $\Gamma$ in the contract form (2.11). We finally obtain that the problems introduced throughout this paper are all equivalent under our standing Assumptions A and B, which is graphically summarised by Figure 1.
Problem 1Problem 2Problem 3Problem 4Lemma 2.5, or [9]Lemma 3.4Theorem 3.9By 2BSDEs in [9]Theorem 3.10Proposition 3.14
@@
We nevertheless remark that, at the end, and depending on the framework under consideration, Problem 2 may still be easier to solve than Problem 4. Indeed, as illustrated through Section 4, the agent’s problem Problem 4 involves a constrained optimisation, making it potentially more complex to solve in practice. However, this does not undermine the effectiveness of our approach, since its main goal is to provide a theoretically simpler method to establish the optimality of the relevant contract form. Ultimately, since all problems are equivalent, this approach offers the flexibility to choose the most practical problem to solve. We also emphasise that the contract form (2.11), originally introduced in [9], is essential for proving the equivalence between the original problem and its ‘first-best’ reformulation. Nevertheless, the key message of our approach is that we do not need to rely on 2BSDEs to demonstrate the optimality of these contracts; their optimality naturally comes from the fact that they allow the principal to achieve the best utility possible.
We believe that this alternative, simpler approach will provide a more straightforward pathway for extending principal–agent problems to more complex frameworks, involving for example multi-agent settings, mean-field interactions, or non-continuous output processes. For instance, solving principal–multi-agent problems with volatility control via the original approach in [9] requires the consideration of ‘multidimensional 2BSDEs’, as highlighted in [18], for which the theory is actually not yet developed. In contrast, following our proposed approach, one could rely on the more mature theory of multidimensional BSDEs to show the equivalence between the original problem and its ‘first-best’ reformulation, thereby ensuring optimality of appropriate revealing contracts without relying on multidimensional 2BSDEs. Furthermore, this simplified approach shall facilitate practical and more widespread implementation of principal–agent problems with volatility control in diverse fields, beyond mathematical finance.
To conclude, we want to emphasise that while our approach provides a simpler and more accessible framework for solving principal–agent problems with volatility control, it relies on two key assumptions: Assumption A, which guarantees weak uniqueness and thus enables the use of the martingale representation theorem, and Assumption B, under which the contract form proposed in [9] remains optimal. First, we view Assumption A as relatively mild and well motivated, particularly given the intended applications. Indeed, without weak uniqueness, the same choice of effort by the agent could generate different distributions for the controlled output process, making the model less meaningful from an economic standpoint. Moreover, although this assumption is central to our current BSDE-based arguments, it may be possible to relax it by considering generalised BSDEs, i.e. with an orthogonal martingale component. This approach would still remain simpler than the general 2BSDE framework, and could be extended for example to multi-agent settings, thanks to the existence of multi-dimensional versions of such BSDEs. Second, we have strong reason to believe that our main result, Theorem 3.10, will remain valid even without Assumption B. However, since the contract form from [9] may no longer be optimal in this case, this would require constructing a new class of contracts. While preliminary exploration in this direction is promising, a full rigorous treatment is left for future research.